Based in Winnipeg, Manitoba, Clearview Security Technologies Inc. (Clearview) was founded to provide security systems, facilities controls, and related services. Clearview established a solid reputation for quality and the business grew, thanks to strong relationships with large long-term customers in Canada and the United States. Clearview has experienced little competitive pressure in its core market and the company's offerings are standardized, enabled by significant technological and financial barriers to entry.

The Research and Innovation Group (RIG) is the development side of the company. Where Clearview's primary lines are standardized, the RIG is all over the map. Clearview uses this smaller division to provide contract software and consulting to a wide range of business types.

The RIG is considering a new contract that will strain resources for not only the RIG, but the entire company. The project involves new technology, a new customer, and a new geographic area. The director of operations has warned you that it will be substantially more risky than anything Clearview does in its core business. With an upfront cost of C$8.5 million, managers want to develop an understanding of expected financing costs. The director of finance explained that understanding cost of capital will be a key part of maintaining and improving Clearview's competitive edge. RIG managers have noticed competing bids for the contract and it is expected that margins will be pushed down.

You have been asked to calculate the company's weighted average cost of capital (WACC), based on the following information. Over the last five years the annual dividends on the firm's stock have grown at 6 percent per year and this growth is expected to continue indefinitely. A dividend of $1.25 was recently paid. Common shares trade at $45 with 250,000 outstanding and no preferred shares. The yield on long-term government bonds is currently 3 percent and you believe the appropriate expected market risk premium is 5 percent (the long-term average). The stock's beta is 1.05. Clearview also has 25-year bonds with $1,000 face value, 6.5-percent semi-annual coupons, and 20 years to maturity. The bonds trade at 94.5. The initial bond offering raised $15.5 million and initially sold at par. The firm's marginal tax rate is 27 percent.

  1. The cost of equity and debt
    1. Calculate Clearview's cost of equity using the constant growth model approach and the CAPM approach. Take the arithmetic average of the two results.
    2. Determine Clearview's after-tax cost of debt. Solving for the cost of debt is best done with a financial calculator, although trial and error will also yield the correct result. The bond valuation formula is required in this approach, substituting values for kb until the bond value is determined. A third method for determining the cost of debt is an approximation formula: kb approx. = ((annual coupon + (face – price)) ÷ years remaining) ÷ ((face + price) ÷ 2). The trial and error737approach can be completed much faster if the approximation method is used first to narrow the trial and error range.
  2. The weight of equity and debt
    1. Calculate the weights of equity and debt in Clearview's capital structure.
    2. Determine Clearview's WACC.
  3. The company will use its current capital structure to set target weights for debt and equity, with flotation costs of 2 percent for long-term debt and 7.5 percent for equity. How much capital must Clearview raise in order to cover the project cost and all flotation charges?
Answer #1

(constant Dividend discount model)

Cost of Equity = (D/P) + g = 1.25/45 + 0.06 = 8.78%

(CAPM model)

Cost of Equity = 3 + 1.05 *5 = 8.25%

Cost of Equity = Avg of above 2 = 8.52%

------------------------------------------------------------------------

Cost of Debt or YTM

PV = 94.5

N = 20 Years

FV = 100

Coupon = 6.5% semi-annual

therefore YTM = 7.01% (Formula or financial calculator)

---------------------------------------------------------------------------

Equity = 45* 250,000 = $11.25 Million

Debt = $ 15.5 Million

Total = $ 26.75 Million

Weight of Debt = 0.5794

Weight of Equity = 0.4205

--------------------------------------------------------------------------

WACC = 0.5794*7.01* (1-0.27) + 0.4205*8.52

= 2.964 + 3.582 = 6.546% (Ans)